OpenStudy (anonymous):

at what vale of x is the line tangent to the graph of y=x^2+3x+5 perpendicular to the line x-2y=5

OpenStudy (bahrom7893):

find the derivative: y = 2x + 3 Slope is 2 in second case slope is: -2y = -x+5 y = 1/2 x- 5/2

OpenStudy (bahrom7893):

oh wait no...

OpenStudy (bahrom7893):

in second case slope is 1/2, so: 1/2 = 2x + 3

OpenStudy (bahrom7893):

1 = 4x + 6 -5 = 4x x = -5/4, check my arithmetic and answer.. but that was right lol

OpenStudy (bahrom7893):

i just didnt finish it..

OpenStudy (anonymous):

that is not a choice. the choices are -5/2, -2, -1/2, 1/2, and 5/2

OpenStudy (bahrom7893):

oh sorry its not supposed to be 1/2 = 2x + 3 because we are looking for perpendicular so: -2 = 2x + 3 -5 = 2x x = -5/2

OpenStudy (anonymous):

alright ty again ignore this it just helps me to type it out. find the slope of first line and plug it into the y for the second line and solve for x